Denton Company manufactures and sells a single product. Cost data for the product are given:

Variable costs per unit:
Direct materials $ 5
Direct labor 11
Variable manufacturing overhead 4
Variable selling and administrative 1
Total variable cost per unit $ 21
Fixed costs per month:
Fixed manufacturing overhead $ 90,000
Fixed selling and administrative 169,000
Total fixed cost per month $ 259,000

The product sells for $54 per unit. Production and sales data for July and August, the first two months of operations, follow:

Units
Produced
Units
Sold
July 30,000 26,000
August 30,000 34,000

The company’s Accounting Department has prepared the following absorption costing income statements for July and August:

July August
Sales $ 1,404,000 $ 1,836,000
Cost of goods sold 598,000 782,000
Gross margin 806,000 1,054,000
Selling and administrative expenses 195,000 203,000
Net operating income $ 611,000 $ 851,000

Required:

1. Determine the unit product cost under:

a. Absorption costing.

b. Variable costing.

2. Prepare variable costing income statements for July and August.

3. Reconcile the variable costing and absorption costing net operating incomes.
Answer #1
1a
Direct materials 5
Direct labor 11
Variable manufacturing overhead 4
Fixed manufacturing overhead 3 =90000/30000
Absorption costing unit product cost 23
1b
Direct materials 5
Direct labor 11
Variable manufacturing overhead 4
Variable costing unit product cost 20
2
July August
Sales 1404000 1836000
Variable expenses:
Variable cost of goods sold 520000 680000
Variable selling and administrative expenses 26000 34000
Total Variable expenses 546000 714000
Contribution margin 858000 1122000
Fixed expenses:
Fixed manufacturing overhead 90000 90000
Fixed selling and administrative expenses 169000 169000
Total Fixed expenses 259000 259000
Net operating income(loss) 599000 863000
3
July August
Variable costing net operating income 599000 863000
Add(deduct) fixed manufacturing overhead deferred in
(released from) inventory under absorption costing
12000 (12000) =4000*3
Absorption costing net operating income 611000 851000
Workings:
July August
Sales =26000*54 =34000*54
Variable expenses:
Variable cost of goods sold =26000*20 =34000*20
Variable selling and administrative expenses =26000*1 =34000*1